# CoPRS: Learning Positional Prior from Chain-of-Thought for Reasoning Segmentation

### v1

@@ Abstract
Existing works on reasoning segmentation either connect hidden features from a language model directly to a mask decoder or represent positions in text, which limits interpretability and semantic detail. To solve this, we present CoPRS, a Multi-modal Chain-of-Thought (MCoT)–based positional perception model that bridges language reasoning to segmentation through a differentiable and interpretable positional prior instantiated as a heatmap. By making the reasoning process clear via MCoT and expressing it as a dense, differentiable heatmap, this interface enhances interpretability and diagnostic analysis and yields more concentrated evidence on the target. A learnable concentration token aggregates features of the image and reasoning text to generate this positional prior, which is decoded to precise masks through a lightweight decoder, providing a direct connection between reasoning and segmentation. Across the RefCOCO series and ReasonSeg, CoPRS matches or surpasses the best reported metrics on each standard split under comparable protocols, with performance at or above prior state of the art across both validation and test partitions. Extensive experiments reveal that the quality of the heatmap strongly influences the resulting mask quality, supporting a consistent association between the reasoning output and downstream mask generation. Collectively, these findings support the utility of this paradigm in bridging reasoning and segmentation and show advantages in concentration driven by reasoning and predicting masks more precisely. Code, checkpoints and logs are released at https://github.com/ZhenyuLU-Heliodore/CoPRS.git.
@@ 1 Introduction
Visual perception is increasingly expected to not only assign labels to pixels but also follow natural-language instructions with compositional constraints, such as “Segment the UAV trailing the quadcopter and is partially occluded by trees.” This demand advances the long arc of visual understanding, starting from semantic segmentation (category labels) (Guo et al., 2018), to instance segmentation (object masks) (Hafiz & Bhat, 2020), and further to open-vocabulary segmentation (open-set text categories) (Ren et al., 2024a), and most recently, toward reasoning segmentation (free-form instructions) Lai et al. (2024). Meeting this goal requires coupling language reasoning with spatial grounding by converting textual instructions into perceptual decisions.
@@
To close this gap, we introduce CoPRS, a Multi-modal Chain-of-Thought (MCoT)–based positional perception model for reasoning segmentation. CoPRS is one-stage and end-to-end: given an image–instruction input, it first reasons before producing a perception heatmap concentrating the target region, which provides a positional prior to enhance the segmentation mask decoding. As compared in Figure 1, the positional prior serves as a differentiable and interpretable connection between MCoT (Wang et al., 2025) and segmentation, which is direct and effective to enhance visual perception of Multi-modal Large Language Model (MLLM) and align instruction semantics with mask decoding.
@@
CoPRS matches or exceeds the best reported cIoU/gIoU on each split under comparable protocols across RefCOCO, RefCOCO+ (Kazemzadeh et al., 2014), RefCOCOg (Mao et al., 2016), and ReasonSeg (Lai et al., 2024). We further find that the quality of the heatmap strongly influences the final mask accuracy. Overall, the results show strong concentration driven by reasoning and precise mask generation. Beyond reasoning segmentation, the unified framework and its positional prior naturally extend to region concentration tasks such as referring tracking and trajectory prediction.
@@
  • Strong Results. CoPRS performs strongly on each split across the RefCOCO series and ReasonSeg, and further analysis clarifies how reasoning output affects segmentation performance.
@@ 2 Related Work
Referring and Reasoning Segmentation. Referring segmentation requires a model to produce a mask for the entity described in a short instruction. Prior work such as VLT (Ding et al., 2021), CRIS (Wang et al., 2022), LAVT (Yang et al., 2022), ReLA (Liu et al., 2023a), X-Decoder (Zou et al., 2023a), SEEM (Zou et al., 2023b), Grounded-SAM (Ren et al., 2024a), typically relies on specific text encoders rather than large language models (LLMs) to parse the text and predict the mask. Reasoning segmentation extends this setting to longer, compositional instructions with stricter grounding requirements, motivating the two method families outlined next.
@@ 3 Method
We first presents the model design and data flow in Section 3.1. We then formalize the learning objectives, unifying policy optimization via GRPO on the language path with segmentation supervision on the vision path in Section 3.2. Finally, we detail the training and inference procedures in Section 3.3, including data preparation, tokenization, group rollouts, and the deterministic inference.
@@ 3.1 Model Architecture
Figure 2: Overall architecture. Given image and text inputs, the policy generates CoT and concentration tokens, which query image keys to generate positional prior, and then decoded to masks. The policy and segmentation modules are jointly optimized.
@@
where $\bm{Q}\!\in\!\mathbb{R}^{d_{q}}$, $\bm{K}\!\in\!\mathbb{R}^{H\times W\times d_{k}}$, $\bm{W}^{Q}_{i}\!\in\!\mathbb{R}^{d_{q}\times d_{h}}$, $\bm{W}^{K}_{i}\!\in\!\mathbb{R}^{d_{k}\times d_{h}}$, $d_{h}$ is the head dimension, $n_{\text{head}}$ is the number of heads, and $\mathcal{F}_{\text{fuse}}:\mathbb{R}^{n_{\text{head}}\times H\times W}\!\to\!\mathbb{R}^{H\times W}$.
Lightweight Decoder. Our mask decoder comprises two submodules. First, three stacked 2D convolutional blocks resample the fused positional prior, producing a feature map at the decoder resolution. Second, we choose a Two-Way Transformer following the SAM decoder desigh (Kirillov et al., 2023): which performs bidirectional cross attention between the image features and the positional prior. This lightweight design has 4.7M parameters and enables the prior to guide dense segmentation. Formally, we formulate the process as
@@ 3.2 Learning Objectives
Unified Objective. We train the whole system end-to-end with a single objective that couples reinforcement learning on the language path with segmentation supervision on the vision path. For each $\left(\bm{x}_{\text{img}},\,\bm{x}_{\text{txt}}\right)$, the policy $\pi_{\bm{\theta}}$ rolls out a group of responses $\bigl\{\bm{y}_{\scriptscriptstyle 1:T_{i}}^{(i)}\bigr\}_{i=1}^{G}$ with the group size $G$, and we compute a GRPO loss $\mathcal{L}_{\text{{grpo}}}$ from the advantages. In parallel, the positioanl prior $\bm{H}_{\text{prior}}$ and the predicted mask $\hat{\bm{M}}$ are supervised against the ground truth mask $\bm{M}_{\text{gt}}$ to yield the segmentation loss $\mathcal{L}_{\text{{seg}}}$. The overall objective is
@@
We compute both terms on the same batch and take a single backward pass through all trainable modules.
@@
For each sampled response in the group, we design a reward function that combines mask quality and CoT format compliance. In Specifically, the mask reward score aggregates soft IoU, soft dice, and hard IoU, while the CoT format reward score is computed via multiple regular expressions for the string matching. We then normalize both rewards to the range $[0,1]$ using fixed coefficients. Further implementation details are provided in Section 4.1.
Supervised Segmentation Objective. The segmentation loss comprises three complementary terms. (i) A binary cross entropy (BCE) loss applied to $\bm{H}_{\text{prior}}$ encourages positional evidence and with accurate concentration. (ii) A dice loss (Milletari et al., 2016) on the predicted mask $\hat{\bm{M}}$ directly supervises mask quality. (iii) A focal loss (Lin et al., 2017) on the mask logits emphasizes hard pixels and fine-grained structures. All losses are computed only over the original image region and averaged per image over the batch, with the dice loss coefficient $\lambda_{d}$ and focal loss coefficient $\lambda_{f}$ being reported in Section 4.1. Formally, the segmentation loss is
@@ 3.3 Training and Inference
Training Procedure. As shown in Figure 2, during training procedure we tokenize $\left(\bm{x}_{\text{img}},\,\bm{x}_{\text{txt}}\right)$, replicate each pair for $G$ times, and feed these copies to the $\pi_{\bm{\theta}}$ to generate $G$ responses. For each response in the group, the reward function assigns a scalar score, and the scores are converted into advantages for computing $\mathcal{L}_{\text{{grpo}}}$, which updates only the MLLM parameters. In the same batch, $\bm{x}_{\text{img}}$ is resized and padded encoded by the vision backbone, and decoded to $\hat{\bm{M}}$ for computing $\mathcal{L}_{\text{{seg}}}$, which updates all trainable modules. We optimize both losses jointly in each iteration.
@@ 4 Experiments
Research Questions. We evaluate the CoPRS by conducting experiments on four datasets. We aim to answer the following research questions:
@@
    Is the positional prior accurately reflect concentration and reasoning, i.e., does its quality strongly correlates with the final segmentation accuracy?
@@
    Do both the GRPO term and the supervised segmentation loss terms contribute to performance, and does the unified objective outperform either component alone?
@@ 4.1 Experimental Setup
Datasets and Metrics. We train CoPRS-3B and CoPRS-7B separately on the training sets of RefCOCO, RefCOCO+ and RefCOCOg. To prevent data leakage, we remove from the training data all COCO images that appear in the validation or test splits of RefCOCO(+/g). We evaluate on the official validation and test splits of RefCOCO(+/g). We further assess zero-shot reasoning segmentation by evaluating on ReasonSeg (validation and test) without training on its images. Consistent with common practice in prior work (e.g. Lai et al. (2024)), we adopt intersection over union (IoU) metrics. Specifically, we report cIoU (the cumulative intersection over the cumulative union) on RefCOCO(+/g), and both cIoU and gIoU (mean of per-image IoU) on ReasonSeg.
@@ Methods
Baselines. We compare our method with 20 prior works with grouping them into three categories. Methods without LLMs, including VLT (Ding et al., 2021), CRIS (Wang et al., 2022), LAVT (Yang et al., 2022), ReLA (Liu et al., 2023a), X-Decoder (Zou et al., 2023a), SEEM (Zou et al., 2023b), Grounded-SAM (Ren et al., 2024a), do not rely LLM to encode instruction texts for generating masks. Latent reasoning methods, including LISA (Lai et al., 2024), PerceptionGPT (Pi et al., 2024), PixelLM (Ren et al., 2024b), LaSagnA (Wei et al., 2024), SegLLM (Ren et al., 2024b), OMG-LLaVA (Zhang et al., 2024a), GroundHog (Zhang et al., 2024b), GLaMM (Rasheed et al., 2024), RAS (Cao et al., 2025), take hidden features from a large language model and decode them into segmentation masks. Text-based reasoning methods, including SAM4MLLM (Chen et al., 2024), Seg-Zero (Liu et al., 2025), Seg-R1 (You & Wu, 2025), Text4Seg (Lan et al., 2025), use a MLLM to emit discrete location tokens—box/point coordinates or patch indices, and then convert them to masks. For approaches available in multiple parameter scales, we report results for all the variants. RAS provides only a version with 13B parameters.
@@ 4.2 Overall Performance (RQ1)
We compare our CoPRS with prior state-of-the-art reasoning segmentation methods on two standard benchmarks: the RefCOCO series and ReasonSeg.
@@ Methods
Results on ReasonSeg. We evaluate on the ReasonSeg in a zero-shot setting to validate the generalization ability of CoPRS on complex reasoning segmentation scenarios. From Table 2, our CoPRS also demonstrates superior results on the complex reasoning segmentation task. Meanwhile, we find that methods trained with reinforcement learning, such as Seg-R1, Seg-Zero and our CoPRS, consistently outperform other methods, demonstrating the generalization benefits of reinforcement learning for segmentation models.
@@ 4.3 Correlation Analysis and Visualization (RQ2)
Correlation Analysis. We analyze the correlation between the positional prior $\bm{H}_{\text{prior}}$ and the predicted mask $\hat{\bm{M}}$ during both training and inference. We plot the corresponding training losses and evaluation metrics as scatter points to make the relationship clear. Additionally, we use ordinary least square regression to plot the regression line $y=\hat{\alpha}+\hat{\beta}x$ and the mean confidence bands $\hat{y}(x)\pm\eta\,\mathrm{s.e.}\left(\hat{y}(x)\right)$, where $\mathrm{s.e.}(\hat{y})=\hat{\sigma}\sqrt{\tfrac{1}{n}+\tfrac{(x-\bar{x})^{2}}{\sum_{i}(x_{i}-\bar{x})^{2}}}$ with $\eta=10$ for visual clarity and $\hat{\sigma}$ being residual standard error.
@@
During Training. Panels (a)–(d) in Figure 3 show blue points, each representing one training batch. The x-axis is $1-\mathcal{L}_{\text{{bce}}}\bigl(\bm{H}_{\text{prior}},\,\bm{M}_{\text{gt}}\bigr)$, which increases as the prior better matches $\bm{M}_{\text{gt}}$. The y-axis is $1-\mathcal{L}_{\text{{dice}}}\bigl(\hat{\bm{M}},\,\bm{M}_{\text{gt}}\bigr)$, which is higher when $\hat{\bm{M}}$ converges to $\bm{M}_{\text{gt}}$. The points exhibit low dispersion, reflecting stable loss with a batch size of 128. Across all datasets, the scatter patterns and correlation coefficients $R>0.7$ indicate a strong positive association between $\bm{H}_{\text{prior}}$ and $\hat{\bm{M}}$.
During Inference. Panels (e)–(h) in Figure 3 show red points, each representing one inference instance. The x-axis is IoU between $\bm{H}_{\text{prior}}$ and $\bm{M}_{\text{gt}}$, i.e., the mask quality if the prior were used directly with no decoding. The y-axis is the IoU between $\hat{\bm{M}}$ and $\bm{M}_{\text{gt}}$, a standard segmentation metric. As in training, the scatter pattern and correlations $R>0.7$ reveal a strong positive relationship across test splits. Is is noticed that the regression lines, confidence bands and most points lie above $y=x$. This trend indicates that the positional already concentrates well, while the decoder further refines it to a precise mask.
Visualization Results. We present zero-shot visualizations on ReasonSeg, as shown in Figure 4. After MCoT reasoning, the positional prior indicates all instances relevant to the instruction (yellow), with the target instance most strongly concentrated (deep red). Figures 6 and 5 present additional successful cases and failure cases (see Appendix Appendix C).

### v2

@@ Abstract
Existing works on reasoning segmentation either connect hidden features from a language model directly to a mask decoder or represent positions in text, which limits interpretability and semantic detail. To solve this, we present CoPRS, a Multi-modal Chain-of-Thought (MCoT)–based positional perception model that bridges language reasoning to segmentation through a differentiable and interpretable positional prior instantiated as a heatmap. By making the reasoning process clear via MCoT and expressing it as a dense, differentiable heatmap, this interface enhances interpretability and diagnostic analysis and yields more concentrated evidence on the target. A learnable concentration token aggregates features of the image and reasoning text to generate this positional prior, which is decoded to precise masks through a lightweight decoder, providing a direct connection between reasoning and segmentation. Across the RefCOCO series and ReasonSeg, CoPRS matches or surpasses the best reported metrics on each standard split under comparable protocols, with performance at or above the prior state of the art across both validation and test partitions. Extensive experiments demonstrate a strong positive correlation among the CoT trajectory, the generated heatmap, and the decoded mask, supporting an interpretable alignment between the reasoning output and downstream mask generation. Collectively, these findings support the utility of this paradigm in bridging reasoning and segmentation and show advantages in concentration driven by reasoning and in more precise mask prediction. Code, checkpoints and logs will be released.
@@ 1 Introduction
Visual perception is increasingly expected to not only assign labels to pixels but also follow natural-language instructions with compositional constraints, such as “Segment the UAV that is trailing the quadcopter and partially occluded by trees.” This demand advances the long arc of visual understanding, starting from semantic segmentation (category labels) (Guo et al., 2018), to instance segmentation (object masks) (Hafiz & Bhat, 2020), and further to open-vocabulary segmentation (open-set text categories) (Ren et al., 2024a), and most recently, toward reasoning segmentation (free-form instructions) Lai et al. (2024). Meeting this goal requires coupling language reasoning with spatial grounding by converting textual instructions into perceptual decisions.
@@
To close this gap, we introduce CoPRS, a CoT-based Positional perception model for Reasoning Segmentation. CoPRS is one-stage and end-to-end: given an image–instruction input, it first reasons before producing a perception heatmap concentrating the target region, which provides a positional prior to enhance the segmentation mask decoding. As compared in Figure 1, the positional prior serves as a differentiable and interpretable connection between MCoT (Wang et al., 2025b) and segmentation, which is direct and effective to enhance visual perception of a Multi-modal Large Language Model (MLLM) and align instruction semantics with mask decoding.
@@
CoPRS matches or exceeds the best reported cIoU/gIoU on each split under comparable protocols across RefCOCO, RefCOCO+ (Kazemzadeh et al., 2014), RefCOCOg (Mao et al., 2016), and ReasonSeg (Lai et al., 2024). We further find a strong positive correlation among the quality of the CoT trajectory, the generated heatmap, and the decoded mask, indicating strong concentration driven by reasoning and precise mask generation. Beyond reasoning segmentation, the unified framework and its positional prior naturally extend to region concentration tasks such as referring tracking and trajectory prediction.
@@
  • Strong Results. CoPRS performs strongly on each split across the RefCOCO series and ReasonSeg, and further analysis clarifies how reasoning output aligns with segmentation performance.
@@ 2 Related Work
Referring and Reasoning Segmentation. Referring segmentation requires a model to produce a mask for the entity described in a short instruction. Prior methods such as VLT (Ding et al., 2021), CRIS (Wang et al., 2022), LAVT (Yang et al., 2022), ReLA (Liu et al., 2023a), X-Decoder (Zou et al., 2023a), SEEM (Zou et al., 2023b), Grounded-SAM (Ren et al., 2024a), typically rely on specific text encoders rather than large language models (LLMs) to parse the text and predict the mask. Reasoning segmentation extends this setting to longer, compositional instructions with stricter grounding requirements, motivating the two method families outlined next.
@@ 3 Method
We first present the model design and data flow in Section 3.1. We then formalize the learning objectives, unifying policy optimization via GRPO on the language path with segmentation supervision on the vision path in Section 3.2. Finally, we detail the training and inference procedures in Section 3.3, including data preparation, tokenization, group rollouts, and deterministic inference.
@@ 3.1 Model Architecture
Figure 2: Overall architecture. Given image and text inputs, the policy generates CoT and concentration tokens, which query image keys to generate a positional prior, that is then decoded to masks. The policy and segmentation modules are jointly optimized.
@@
where $\bm{Q}\!\in\!\mathbb{R}^{d_{q}}$, $\bm{K}\!\in\!\mathbb{R}^{H\times W\times d_{k}}$, $\bm{W}^{Q}_{i}\!\in\!\mathbb{R}^{d_{q}\times d_{h}}$, $\bm{W}^{K}_{i}\!\in\!\mathbb{R}^{d_{k}\times d_{h}}$, $d_{h}$ is the head dimension, $n_{\text{head}}$ is the number of heads, and $\mathcal{F}_{\text{fuse}}:\mathbb{R}^{n_{\text{head}}\times H\times W}\!\to\!\mathbb{R}^{H\times W}$. Details are provided in Algorithm 1.
Lightweight Decoder. Our mask decoder comprises two submodules. First, three stacked 2D convolutional blocks resample the fused positional prior, producing a feature map at the decoder resolution. Second, we choose a Two-Way Transformer following the SAM decoder design (Kirillov et al., 2023), which performs bidirectional cross attention between the image features and the positional prior. This lightweight design has 4.7M parameters and enables the prior to guide dense segmentation. Formally, we formulate the process as
@@ 3.2 Learning Objectives
Unified Objective. We train the whole system end-to-end with a single objective that couples reinforcement learning on the language path with segmentation supervision on the vision path. For each $\left(\bm{x}_{\text{img}},\,\bm{x}_{\text{txt}}\right)$, the policy $\pi_{\bm{\theta}}$ rolls out a group of responses $\bigl\{\bm{y}_{\scriptscriptstyle 1:T_{i}}^{(i)}\bigr\}_{i=1}^{G}$ with the group size $G$, and we compute a GRPO loss $\mathcal{L}_{\text{{grpo}}}$ from the advantages. In parallel, the positional prior $\bm{H}_{\text{prior}}$ and the predicted mask $\hat{\bm{M}}$ are supervised against the ground truth mask $\bm{M}_{\text{gt}}$ to yield the segmentation loss $\mathcal{L}_{\text{{seg}}}$. The overall objective is
@@
We compute both terms for each batch and take a single backward pass through all trainable modules.
@@
For each sampled response in the group, we design a reward function that combines mask quality and CoT format compliance. Specifically, the mask reward score aggregates soft IoU, soft dice, and hard IoU, while the CoT format reward score is computed via multiple regular expressions for the string matching. We then normalize both rewards to the range $[0,1]$ using fixed coefficients. Further implementation details are provided in Section 4.1.
Supervised Segmentation Objective. The segmentation loss comprises three complementary terms. (i) A binary cross-entropy (BCE) loss applied to $\bm{H}_{\text{prior}}$ encourages positional evidence and accurate concentration. (ii) A dice loss (Milletari et al., 2016) on the predicted mask $\hat{\bm{M}}$ directly supervises mask quality. (iii) A focal loss (Lin et al., 2017) on the mask logits emphasizes hard pixels and fine-grained structures. All losses are computed only over the original image region and averaged per image over the batch, with the dice loss coefficient $\lambda_{d}$ and focal loss coefficient $\lambda_{f}$ being reported in Section 4.1. Formally, the segmentation loss is
@@ 3.3 Training and Inference
Training Procedure. As shown in Figure 2, during training we tokenize $\left(\bm{x}_{\text{img}},\,\bm{x}_{\text{txt}}\right)$, replicate each pair for $G$ times, and feed these copies to the $\pi_{\bm{\theta}}$ to generate $G$ responses. For each response in the group, the reward function assigns a scalar score, and the scores are converted into advantages for computing $\mathcal{L}_{\text{{grpo}}}$, which updates only the MLLM parameters. In the same batch, $\bm{x}_{\text{img}}$ is resized and padded, then encoded by the vision backbone, and decoded to $\hat{\bm{M}}$ for computing $\mathcal{L}_{\text{{seg}}}$, which updates all trainable modules. We optimize both losses jointly in each iteration.
@@ 4 Experiments
Research Questions. In this section, we aim to answer the following research questions:
@@
    How are the CoT, the positional prior $\bm{H}_{\text{prior}}$, and the predicted mask $\hat{\bm{M}}$ mutually correlated, i.e., does higher CoT quality align with stronger positional priors and better segmentation accuracy?
@@
    Do the GRPO settings, supervised segmentation losses, and MLLM/vision backbone choices each contribute to performance, and does our unified objective with the default backbones outperform these alternatives?
@@ 4.1 Experimental Setup
Datasets and Metrics. We evaluate CoPRS by conducting experiments on four datasets. We train CoPRS-3B and CoPRS-7B separately on the training sets of RefCOCO, RefCOCO+ and RefCOCOg. To prevent data leakage, we remove from the training data all COCO images that appear in the validation or test splits of RefCOCO(+/g). We evaluate on the official validation and test splits of RefCOCO(+/g). We further assess zero-shot reasoning segmentation by evaluating on ReasonSeg (validation and test) without training on its images. Consistent with common practice in prior work (e.g., Lai et al. (2024)), we adopt intersection over union (IoU) metrics. Specifically, we report cIoU (the cumulative intersection over the cumulative union) on RefCOCO(+/g), and both cIoU and gIoU (mean of per-image IoU) on ReasonSeg.
@@ Methods
Baselines. We compare our method with 20 prior works grouped into three categories. Methods without LLMs, including VLT (Ding et al., 2021), CRIS (Wang et al., 2022), LAVT (Yang et al., 2022), ReLA (Liu et al., 2023a), X-Decoder (Zou et al., 2023a), SEEM (Zou et al., 2023b), Grounded-SAM (Ren et al., 2024a), do not rely on LLM to encode instruction texts for generating masks. Latent reasoning methods, including LISA (Lai et al., 2024), PerceptionGPT (Pi et al., 2024), PixelLM (Ren et al., 2024b), LaSagnA (Wei et al., 2024), SegLLM (Wang et al., 2025a), OMG-LLaVA (Zhang et al., 2024a), GroundHog (Zhang et al., 2024b), GLaMM (Rasheed et al., 2024), RAS (Cao et al., 2025), take hidden features from a large language model and decode them into segmentation masks. Text-based reasoning methods, including SAM4MLLM (Chen et al., 2024), Seg-Zero (Liu et al., 2025), Seg-R1 (You & Wu, 2025), Text4Seg (Lan et al., 2025), use an MLLM to emit discrete location tokens—box/point coordinates or patch indices, and then convert them to masks. For approaches available in multiple parameter scales, we report results for all the variants. RAS provides only a version with 13B parameters.
@@ 4.2 Overall Performance (RQ1)
We compare CoPRS with prior state-of-the-art reasoning segmentation methods on two standard benchmarks: the RefCOCO series and ReasonSeg.
@@ Methods
Results on ReasonSeg. We evaluate on ReasonSeg in a zero-shot setting to validate the generalization ability of CoPRS on complex reasoning segmentation scenarios. From Table 2, our CoPRS also demonstrates superior results on the complex reasoning segmentation task. Meanwhile, we find that methods trained with reinforcement learning, such as Seg-R1, Seg-Zero and our CoPRS, consistently outperform other methods, demonstrating the generalization benefits of reinforcement learning for segmentation models.
@@ 4.3 Correlation Analysis and Visualization (RQ2)
Correlation Analysis Methodology. We first analyze the correlation between the positional prior $\bm{H}_{\text{prior}}$ and the predicted mask $\hat{\bm{M}}$ during both training and inference. We then analyze how the quality of CoT correlates with both $\bm{H}_{\text{prior}}$ and $\hat{\bm{M}}$, thereby linking the linguistic reasoning to the visual outputs. We plot the corresponding training losses and evaluation metrics as scatter points to make the relationship clear. Additionally, we use ordinary least square regression to plot the regression line $y=\hat{\alpha}+\hat{\beta}x$ and the mean confidence bands $\hat{y}(x)\pm\eta\,\mathrm{s.e.}\left(\hat{y}(x)\right)$, where $\mathrm{s.e.}(\hat{y})=\hat{\sigma}\sqrt{\tfrac{1}{n}+\tfrac{(x-\bar{x})^{2}}{\sum_{i}(x_{i}-\bar{x})^{2}}}$ with $\eta=10$ for visual clarity and $\hat{\sigma}$ being residual standard error.
@@
Correlation between Heatmap and Mask. During training, panels (a)–(d) in Figure 3 show blue points, each representing one training batch. The x-axis is $1-\mathcal{L}_{\text{{bce}}}\bigl(\bm{H}_{\text{prior}},\,\bm{M}_{\text{gt}}\bigr)$, which increases as the prior better matches $\bm{M}_{\text{gt}}$. The y-axis is $1-\mathcal{L}_{\text{{dice}}}\bigl(\hat{\bm{M}},\,\bm{M}_{\text{gt}}\bigr)$, which is higher when $\hat{\bm{M}}$ converges to $\bm{M}_{\text{gt}}$. The points exhibit low dispersion, reflecting stable loss with batch size of 128. Across all datasets, the scatter patterns and correlation coefficients $R>0.7$ indicate a strong positive association between $\bm{H}_{\text{prior}}$ and $\hat{\bm{M}}$.
Table 3: CoT consistency. Consistency score ranges with sample counts, mean heatmap IoU, and mean mask IoU on RefCOCO+.
Consistency Score #Samples Heatmap mIoU Mask mIoU
$[0,0.25)$ 225 0.25 0.55
$[0.25,0.5)$ 568 0.51 0.78
$[0.5,0.75)$ 749 0.69 0.90
$[0.75,1.0]$ 359 0.82 0.94
Figure 4: Correlation between CoT quality and segmentation quality (Heatmap/Mask IoU) on RefCOCO+. OLS results are overlaid.
During inference, panels (e)–(h) in Figure 3 show red points, each representing one inference instance. The x-axis is IoU between $\bm{H}_{\text{prior}}$ and $\bm{M}_{\text{gt}}$, i.e., the mask quality if the prior were used directly with no decoding. The y-axis is the IoU between $\hat{\bm{M}}$ and $\bm{M}_{\text{gt}}$, a standard segmentation metric. As in training, the scatter pattern and correlations $R>0.7$ reveal a strong positive relationship across test splits. It is observed that the regression lines, confidence bands and most points lie above $y=x$. This trend indicates that the positional prior already concentrates well, while the decoder further refines it to a precise mask.
Correlation between CoT and Segmentation Quality. While Figure 3 already confirms the alignment between the heatmap and the final masks, it does not yet quantify how well the CoT reasoning itself aligns with these visual outputs. To make this link more explicit, we additionally use Gemini-2.5-Flash (Comanici et al., 2025) as an independent automatic evaluator. Inspired by Yin et al. (2025), we compute a consistency score in $[0,1]$ (weighted average over four dimensions: logical correctness 0.3, task relevance 0.2, visual consistency 0.3, localization accuracy 0.2) between the image–instruction pair and the generated CoT on the RefCOCO+ testA split. The scatter plots in Figure 4 show a clear positive correlation between CoT consistency scores and both Heatmap IoU and Mask IoU. Moreover, Section 4.3 groups samples by consistency score range and reports the number of samples and heatmap/mask mIoU in each range. Higher consistency bins consistently achieve higher segmentation quality. This quantitative evidence directly supports that better CoT reasoning quality leads to better segmentation performance in CoPRS.
Visualization Results. We present zero-shot visualizations on ReasonSeg, as shown in Figure 5. After MCoT reasoning, the positional prior indicates all instances relevant to the instruction (yellow), with the target instance most strongly concentrated (deep red). Figure 8 in Appendix presents additional visualizations. Additional failure cases in Figure 7 in the Appendix show that CoPRS mainly struggles with very small objects that disappear at our current input resolution, and dense groups of similar instances where text alone cannot reliably disambiguate the target.